a. A shipment of 7 television sets contains 2 defective sets. A hotel makes a random



purchase of 3 of the sets. If x is the number of defective sets purchased by the hotel,



find the probability distribution of X. Express the results graphically as probability



histogram.



b. Find the cumulative distribution function of the random variable X representing the



number of defective in problem a. Then using F(x). Find



P(X=1) and P (0X2)



c. Construct the cumulative distribution function of problem b

1
Expert's answer
2021-12-17T12:37:01-0500

P(x)=P(7)-P(3)

"P(7)=\frac{7}{3}"

"P(3)=\frac{3}{7}"

"P(x)=\frac{7}{3}-\frac{3}{7}"

="\frac{40}{21}=1.9048"